OpenStudy (anonymous):

At the candy store, a chocolate bar costs $c$ dollars and a vanilla bar costs $2$ dollars more than a chocolate bar. Jamie buys a chocolate bar and three vanilla bars, and Kevin buys five chocolate bars. How much money, in total, do Jamie and Kevin spend at the candy store in terms of $c$?

OpenStudy (texaschic101):

c = chocolate bar v = vanilla bar = c + 2 c + 3(c + 2) is what jamie bought 5c = is what Kevin bought c + 3(c + 2) + 5c = total spent (T) c + 3c + 6 + 5c = T 9c + 6 = T 9c = T - 6 c = (T - 6) / 9 or c = 1/9T - 2/3 you might better get a second opinion on this....